OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given number as solutions. Type the equation in standard form,ax^2+bx+c=0 -√10, 6√10

jimthompson5910 (jim_thompson5910):

\[\large x=-\sqrt{10} \ \ \textrm{or} \ \ x=6\sqrt{10}\] \[\large x+\sqrt{10}=0 \ \ \textrm{or} \ \ x-6\sqrt{10}=0\] \[\large (x+\sqrt{10})(x-6\sqrt{10})=0\] \[\large x^2-6x\sqrt{10}+x\sqrt{10}-60=0\] \[\large x^2-5x\sqrt{10}-60=0\]

jimthompson5910 (jim_thompson5910):

Note: In the answer above, \[\large a=1, \ \ b=-5\sqrt{10}, \ \ \textrm{and} \ \ c=-60\]